OpenStudy (anonymous):

how do i evaluate cos^(-1)(-r2/2)??

OpenStudy (anonymous):

the inverse cosine for -45degrees is undef......?

OpenStudy (anonymous):

question is not clear. looks like \[\cos^{-1}(-\frac{r^2}{2})\]

OpenStudy (anonymous):

it's the inverse cosine of negative root 2 over 2

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

you want a number whose cosine is \[-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

and that number has to be between 0 and pi or 0 and 180 if you are working in degrees

OpenStudy (anonymous):

right i follow the restriction on the range

OpenStudy (anonymous):

degrees or radians?

OpenStudy (anonymous):

just says evaluate. because it's given as -(sqrt2/2) i'm assuming rads

OpenStudy (anonymous):

in any case it will be \[\frac{3\pi}{4}\] or \[135\]

OpenStudy (anonymous):

look at unit circle on last page of trig cheat sheet. you will see that the first coordinate at \[\frac{3\pi}{4}\] is \[-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

so that is your answer

OpenStudy (anonymous):

so the inverse cosine function just tells me the angle associated with the arc length -sqrt2/2 ?